If the numerical value of the area of a retangle is twice the length and the length is 1 more than 3 times the width, what are the dimensions?
area = L * W
w = x
L = 2x + x
2x + x = x(3x + 1

Answer by ewatrrr(24785)


Hi
Let x and (3x + 1)represent the width and length respectively
Might recommend using the above format prior to Writing the Equation
Question states*** A = L*W
L*W = 2*L the question states is true in this question
x(3x+1) = 2(3x + 1)
x = 2, the width. the length is 7 (3*2 + 1)
CHECKING our Answer
2*7 = 2*7